OpenStudy (itrymath):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). (2 points) y = 2x - 2 y = - 1/2x - 7 y = - 1/2 x - 2 y = 2x - 7

OpenStudy (will.h):

parallel means they have the same slope now use y - y1 = m(x - x1) y - 3 = 2(x - 5) y - 3 = 2x - 10 y = 2x - 7

OpenStudy (itrymath):

are for real

OpenStudy (itrymath):

i know this

OpenStudy (itrymath):

omg

OpenStudy (itrymath):

this is what stupid looks like

OpenStudy (will.h):

lol relax you here to learn hope that helps

OpenStudy (itrymath):

it does thanks!

OpenStudy (will.h):

yw!

OpenStudy (itrymath):

will

OpenStudy (itrymath):

what if you changed the 3 to negative 3

OpenStudy (will.h):

y - (-3) = 2(x - 5) y + 3 = 2x - 10 y = 2x -13

OpenStudy (itrymath):

oh and 1 more question.... about this

OpenStudy (will.h):

shoot

OpenStudy (itrymath):

how did you find the slope... where did you get another set of points

OpenStudy (will.h):

When they said " parallel to the line y = 2x + 2 " it means the equation that you will derive has the same slope as the equation above since they are parallel since the equation above has a slope of 2 then the equation that we will derive will have the same slope (2)

OpenStudy (itrymath):

oh

OpenStudy (itrymath):

okay that it thanks!!!!!!!!!!

OpenStudy (itrymath):

i wish i could give more mdeals

OpenStudy (will.h):

parallel lines = same slope perpendicular lines = slopes when multiplied will equal to -1 you welcome happy to help